# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ 5.3 Picard group of ${\rm Sh}(M)$
If $H$ satisfies (5.4), then the map $\Phi(\mathrm{H})\in\mathrm{ACusp}_{\frac{2+n}{2}}(\rho_{M})^{\vee}$ satisfies
$\Phi(\mathrm{H})(f)=0$
for any $f\in\mathrm{Cusp}_{\frac{2+n}{2}}(\rho_{M})$. This means $\Phi(\mathrm{H})$ is proportional to the constant coefficients function $\Phi(\mathrm{H}_{0,0})$. It follows that $\mathrm{H}$ is a multiple of $\mathrm{H}_{0,0}$. The converse is obvious. ∎
6 Heegner divisors on $\overline{{\rm Sh}}(M)$
6.1 Baily-Borel compactification
There is a Baily-Borel compactification $\overline{{\rm Sh}}(M)$ of ${\rm Sh}(M)$ by adding modular curves and cusps, which is a normal projective variety. Abstractly, it is given by the Proj of the graded ring of automorphic forms on $\mathbf{D}$. In the modern language, we have the following isomorphism
$\overline{{\rm Sh}}(M)\cong{\rm Proj}~{}\bigoplus\limits_{m}\mathrm{H}^{0}\left({\rm Sh}(M),\lambda_{M}^{\otimes m}\right).$
In other words, the Hodge line bundle $\lambda_{M}$ can be extended to an ample $\mathbb{Q}$-line bundle on $\overline{{\rm Sh}}(M)$, denoted by $\overline{\lambda}_{M}$.
A more explicit description of $\overline{{\rm Sh}}(M)$ as an arithmetic quotient is given below. If $V\subseteq M$ is a sublattice, we denote by
$\pi_{V}:\mathbb{P}(M\otimes\mathbb{C})-\mathbb{P}(V\otimes{\mathbb{C}})\longrightarrow\mathbb{P}(M/V\otimes\mathbb{C})$
the natural projection. Then we have
$\overline{{\rm Sh}}(M)=\widetilde{\mathrm{O}}(M)\backslash\big{(}\mathbf{D}\bigsqcup\limits_{I:{\rm isotropic~{}line}}\pi_{I^{\perp}}\mathbf{D}\cup\bigsqcup\limits_{J:{\rm isotropic~{}plane}}\pi_{J^{\perp}}\mathbf{D}\big{)}.$
The 1-dimensional boundary components of $\overline{{\rm Sh}}(M)$ correspond one-to-one to the orbits of rank $2$ primitive isotropic sublattices of $M$ under the action of $\widetilde{\mathrm{O}}(M)$. We denote by $\partial_{J}(\overline{{\rm Sh}}(M))\subseteq\partial(\overline{{\rm Sh}}(M))$ the boundary component associated to some isotropic plane $J$.
For each isotropic plane $J\subseteq M$, the lattice $J^{\perp}/J$ is a negative definite lattice of rank $n-2$. A natural question is whether, for any isotropic subgroup $H\subseteq G_{M}$, there is an inclusion
$\mathbf{Gen}_{0,n-2}(H^{\perp}/H)\subseteq\{J^{\perp}/J~{}\mid~{}J\subseteq M~{}\hbox{is an isotropic plane}\}/\cong.$ (6.1)
We provide a sufficient condition for this to be held.
@@ 6.1 Baily-Borel compactification
If $M$ splits two hyperbolic planes, then there is an inclusion
$\mathbf{Gen}_{0,n-2}(G_{M})\subseteq\{J^{\perp}/J~{}\mid~{}J\subseteq M~{}\hbox{is an isotropic plane}\}/\cong.$ (6.2)
@@
For any negative definite lattice $L\in\mathbf{Gen}_{0,n-2}(G_{M})$,
$L\oplus U^{\oplus 2}\in\mathbf{Gen}_{2,n}(G_{M}).$
@@
Since $\sharp~{}|\mathbf{Gen}_{2,n}(G_{M})|=1$ by [27, Theorem 1.13.1] , we have $L\oplus U^{\oplus 2}\cong M$. Now it is sufficient to take $J$ to be the image of the standard isotropic plane of $U^{\oplus 2}$ under this isomorphism.
6.2 Heegner divisors on the boundary
Let $x\in\partial(\overline{{\rm Sh}}(M))$ be a boundary point. The (analytic) local Picard group of $\overline{{\rm Sh}}(M)$ at $x$ is defined by
${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x):=\lim\limits_{x\in U}{\rm Pic}_{\mathbb{Q}}(U\cap{\rm Sh}(M))$
where $U$ runs through all (analytic) open neighbourhoods of $x$. There is a restriction map
We say $x$ is a generic point if it is not a cusp. For any isotropic plane $J\subseteq M$, we say a divisor $H\in{\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ is trivial at generic points of $\partial_{J}(\overline{{\rm Sh}}(M))$ if there exists a generic point $x\in\partial_{J}(\overline{{\rm Sh}}(M))$ such that the image of $H$ in ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x)$ is trivial. The natural projection $\mathrm{O}(M)\rightarrow\mathrm{O}(G_{M})$ is surjective if $M$ splits two hyperbolic planes (cf. [27, Theorem 1.14.2]). Now the result of Bruinier and Fretaig about local Borcherds products immediately implies

### v2

@@ 5.3. Heegner divisors as Coefficient Functionals
If $\mathcal{H}$ satisfies (5.6), then the functional $\Phi(\mathcal{H})\in\mathrm{ACusp}_{\frac{2+n}{2}}(\rho_{M}^{*})^{\vee}$ annihilates all cusp forms, i.e.:
$\Phi(\mathcal{H})(f)=0,\forall f\in\mathrm{Cusp}_{\frac{2+n}{2}}(\rho_{M}^{*}).$
Thus, $\Phi(\mathcal{H})$ is proportional to $\Phi(\mathrm{H}_{0,0})$, the functional that extracts the constant term. Hence, $\mathcal{H}$ is a multiple of $\mathrm{H}_{0,0}$. The converse follows directly. ∎
5.4. Extending Heegner Divisors to the Boundary
We mainly focus on the case ${\rm Sh}(M)$. Let $x\in\partial\overline{{\rm Sh}}(M)$ be a boundary point. The analytic local Picard group of $\overline{{\rm Sh}}(M)$ at $x$ is defined as the direct limit
${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x)\coloneqq\lim_{\begin{subarray}{c}\longrightarrow\\ x\in U\end{subarray}}{\rm Pic}_{\mathbb{Q}}(U\cap{\rm Sh}(M)),$
where $U$ ranges over analytic open neighborhoods of $x$. The natural restriction map
relates global and local Cartier divisor classes.
A point $x\in\partial\overline{{\rm Sh}}(M)$ is called generic if it is not a cusp. Given an isotropic plane $J\subset M$, we say a divisor $H\in{\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ is trivial at generic points of $\partial_{J}\subseteq\overline{{\rm Sh}}(M)$ if there exists a generic point $x\in\partial_{J}$ where the image of $H$ under the restriction map vanishes in ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x)$.
In [BF01], Bruinier and Freitag developed a local analogue of Borcherds products to characterize when a linear combination of Heegner divisors vanishes at generic boundary points. We will restate their criterion in a form adapted to our setting. In the following, we assume the lattice $M$ has the decomposition
$M=U(N_{1})\oplus U(N_{2})\oplus L$ (5.8)
where $L$ is an even definite lattice, $N_{1},N_{2}\in\mathbb{Z}\setminus\{0\}$. This decomposition induces a decomposition of discriminant groups
$G_{M}=G_{U(N_{1})\oplus U(N_{2})}\oplus G_{L}.$
Let $e_{i},f_{i}\in U(N_{i})$ be the standard basis of $U(N_{i})$. Let $J\subset M$ be the standard isotropic plane generated by $e_{1}$ and $e_{2}$, and define the isotropic subgroup $H_{J}\subset G_{M}$ as the span of $e_{1}/N_{1}$ and $e_{2}/N_{2}$. Then the composition
$G_{L}\subset H_{J}^{\perp}\twoheadrightarrow H_{J}^{\bot}/H_{J}$ (5.9)
is an isomorphism. Under this isomorphism, we let $\pi:H_{J}^{\bot}\twoheadrightarrow G_{L}$ denote the projection map induced by $H_{J}^{\perp}\twoheadrightarrow H_{J}^{\bot}/H_{J}$. The following result is a specialization of one direction of [BF01, Thm. 5.1]:
Theorem 5.11.
Let $x\in\partial_{J}$ be a generic point. If a finite linear combination
$\sum_{\gamma\in G_{M}}\sum_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}$
has trivial image in ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x)$, then for every isotropic-lifted theta series
$\uparrow_{H_{J}}(\Theta_{L,F})=\sum_{\gamma\in G_{M}}\sum_{m\in\mathbb{Q}}c_{m,\gamma}q^{m}\mathbf{e}_{\gamma},$
where $F$ ranges over degree 2 harmonic polynomials, we have
$\sum_{\gamma\in G_{M}}\sum_{m\in\mathbb{Q}}a_{m,\gamma}c_{-m,\gamma}=0.$
Proof.
By [BF01, Lem. 4.1] and subsequent remarks, the image of $\mathrm{H}_{m,\gamma}$ in ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x)$ vanishes when $\gamma\notin H_{J}^{\perp}$. Therefore, the triviality condition reduces to
$\sum_{\gamma\in H_{J}^{\perp}}\sum_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}=0\quad\text{in }{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),x).$
Applying [BF01, Thm. 5.1] to this reduced sum, we obtain
$\sum_{\gamma\in H_{J}^{\perp}}\sum_{m\in\mathbb{Q}}a_{m,\gamma}c^{\prime}_{-m,\pi(\gamma)}=0$
for all theta series
$\Theta_{L,F}=\sum_{\gamma\in G_{L}}\sum_{m\in\mathbb{Q}}c^{\prime}_{m,\gamma}q^{m}\mathbf{e}_{\gamma},$
where $F$ runs over degree 2 harmonic polynomials.
By the definition of isotropic lifting, $\uparrow_{H_{J}}(\Theta_{L,F})$ satisfies $c_{m,\gamma}=c^{\prime}_{m,\pi(\gamma)}$ if $\gamma\in H_{J}^{\perp}$, and $c_{m,\gamma}=0$ otherwise. Thus, the above equality extends to
$\sum_{\gamma\in G_{M}}\sum_{m\in\mathbb{Q}}a_{m,\gamma}c_{-m,\gamma}=0,$
completing the proof. ∎
5.5. An obstruction space
Based on Theorem 5.11, we introduce a subspace of ${\mathrm{Cusp}}_{k}(\rho_{M}^{*})$ capturing obstructions to extending Heegner divisor classes to the boundary. Let $M$ be an even lattice admitting an orthogonal decomposition
$M=U(N_{1})\oplus U(N_{2})\oplus L,$ (5.10)
where:
  • $L$ is a negative definite even lattice;
  • $N_{1},N_{2}\in\mathbb{Z}_{\neq 0}$.
We say $J\subset M$ is an admissible isotropic plane, if $J$ is the standard isotropic plane in $U(N_{1})\oplus U(N_{2})\subset M$ for some decomposition of $M$ as (5.10). Let $H_{J}\subseteq G_{M}$ be the associated isotropic subgroup.
Definition 5.12.
The obstruction space is defined as
${\mathrm{Cusp}}^{\mathrm{obs}}_{\frac{n+2}{2}}(\rho_{M}^{*})\coloneqq\mathrm{Span}\left\{\,\uparrow_{H_{J}}(\sigma^{*}\Theta_{J^{\perp}/J,F})\ \bigg{|}\ \begin{aligned} &J\subset M\hbox{~{}admissible},\\ &F~{}\hbox{harmonic of degree 2},\\ &\sigma\in\mathrm{O}(H_{J}^{\bot}/H_{J})\end{aligned}\right\}.$
Let ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))^{\rm Heegner}\subseteq{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))$ be the subspace consisting of elements whose restriction to ${\rm Sh}(M)$ is lying in ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))^{\rm Heegner}$. A key observation is
@@ 5.5. An obstruction space
Let $M$ be an even lattice as above, and further assume $M$ splits a hyperbolic plane. If
${\mathrm{Cusp}}_{\frac{n+2}{2}}^{\rm obs}(\rho_{M}^{\ast})={\mathrm{Cusp}}_{\frac{n+2}{2}}(\rho_{M}^{\ast}),$
then ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))^{\rm Heegner}\cong\mathbb{Q}$ is spanned by the extended Hodge bundle.
@@
As $\overline{{\rm Sh}}(M)$ is normal, one can view ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))^{\rm Heegner}$ as a subspace of ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))^{\rm Heegner}$ via the following commutative diagram:
[…]
Here, the second horizontal map is an isomorphism because the boundary components of ${\rm Sh}(M)$ in $\overline{{\rm Sh}}(M)$ have codimension $>1$.
For any element $\mathcal{H}\in{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))^{\rm Heegner}$, write
$\mathcal{H}=\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}.$
As $\mathcal{H}$ is $\mathbb{Q}$-Cartier, its image in each local Picard group is a torsion element. In particular, for an admissible isotropic plane $J$, $\mathcal{H}$ is trivial at generic points of $\partial_{J}\subseteq\overline{{\rm Sh}}(M)$. Hence by Theorem 5.11, the pairing of $\mathcal{H}$ and $\uparrow_{H_{J}}(\Theta_{L,F})$ vanishes, when $F$ runs over degree 2 harmonic polynomials.
Now suppose an admissible isotropic plane $J$ arises from a decomposition
$M=U(N_{1})\oplus U(N_{2})\oplus L,$
then $G_{M}$ has a decomposition of the form
$G_{M}=G_{U(N_{1})\oplus U(N_{2})}\oplus G_{L}.$
Fix the isomorphism between $G_{L}$ and $H_{J}^{\perp}/H_{J}$ as (5.9). For each $\sigma\in\mathrm{O}(H_{J}^{\bot}/H_{J})$, it can be transferred to an element in $\mathrm{O}(G_{L})$ and can be extended diagonally by identity on $G_{U(N_{1})\oplus U(N_{2})}$ to an element $\sigma^{\prime}\in\mathrm{O}(G_{M})$. By [Ni79, Thm. 1.14.2], the natural map
$\mathrm{O}(M)\rightarrow\mathrm{O}(G_{M})$
is surjective, so $\sigma^{\prime}$ can be further lifted to an element $\tilde{\sigma}\in\mathrm{O}(M)$. The same argument applied to $\tilde{\sigma}(J)$ imples that the pairing of $\mathcal{H}$ and $\uparrow_{H_{J}}(\sigma^{*}\Theta_{L,F})$ vanishes, when $F$ runs over degree 2 harmonic polynomials.
In summary, the pairing of $\mathcal{H}$ with any element in $\mathrm{Cusp}_{\frac{n+2}{2}}^{\mathrm{obs}}(\rho_{M}^{\ast})$ vanishes. Now the equality $\mathrm{Cusp}_{\frac{n+2}{2}}^{\mathrm{obs}}(\rho_{M}^{\ast})=\mathrm{Cusp}_{\frac{n+2}{2}}(\rho_{M}^{\ast})$ combined with Theorem 5.10 forces $\mathcal{H}$ proportional to the Hodge bundle. ∎
6. Picard group of certain Baily–Borel compactifications
In this section, we analyze the negative definite lattices arising in the Baily–Borel compactifications. By applying obstruction theory of extending Heegner divisors, we establish our main theorem. Additionally, we extend the work of [DFV24] to derive a torsion freeness result for orthogonal Shimura varieties of K3 type.
6.1. Lattices associated to Baily–Borel compactifications
For any isotropic plane $J\subset M$, one can associate a negative definite lattice $J^{\perp}/J$ of rank $n-2$. A natural question arises:
Question. Is every genus representative of $G_{J^{\perp}/J}$-type negative definite lattice associated to some isotropic planes of $M$? More generally, for any isotropic subgroup $H\subset G_{M}$, is there an inclusion
$\mathbf{Gen}_{0,n-2}(H^{\perp}/H)\subseteq\left\{J^{\perp}/J\mid J\subset M\text{ isotropic plane}\right\}/{\cong}\,?$ (6.1)
We establish this under hyperbolic splitting conditions.
Proposition 6.1.
Suppose $M\cong U(N_{1})\oplus U(N_{2})\oplus L$ for some $N_{i}\in\mathbb{Z}$ and a negative definite lattice $L$. If $\gcd(N_{1},N_{2})\mid 2$, we have
$\mathbf{Gen}_{0,n-2}(G_{L})\subseteq\{J^{\perp}/J~{}\mid~{}J\subseteq M~{}\hbox{is an isotropic plane}\}/\cong.$
In particular, when $N_{1}=N_{2}=1$, we have
$\mathbf{Gen}_{0,n-2}(G_{M})\subseteq\{J^{\perp}/J~{}\mid~{}J\subseteq M~{}\hbox{is an isotropic plane}\}/\cong.$
Proof.
Let $L^{\prime}\in\mathbf{Gen}_{0,n-2}(G_{L})$ be a negative definite lattice. Then,
$L^{\prime}\oplus U(N_{1})\oplus U(N_{2})\in\mathbf{Gen}_{2,n}(G_{M}).$
By Nikulin’s uniqueness theorem [Ni79, Thm. 1.13.1], the genus $\mathbf{Gen}_{2,n}(G_{M})$ contains exactly one isomorphism class. Thus, we have
$L^{\prime}\oplus U(N_{1})\oplus U(N_{2})\cong M.$
@@ 6.1. Lattices associated to Baily–Borel compactifications
Let $J$ be the image of the standard isotropic plane in $U(N_{1})\oplus U(N_{2})$ under this isomorphism. Then $J^{\perp}/J\cong L^{\prime}$, proving the inclusion. ∎
Remark 6.2.
From the proof, it is clear that the right-hand side of the inclusion (6.1) may be restricted to range only over admissible isotropic planes (see Subsection 5.5).
6.2. Picard Group of $\overline{{\rm Sh}}_{\Gamma}(M)$
We begin by considering lattices that admit sufficient splitting properties. These lattices serve as a natural setting for applying our arithmetic results.
Definition 6.3.
An even lattice $M$ is of K3 type if it satisfies:
  1. (1)
    $M$ splits two hyperbolic planes over $\mathbb{Z}$;
  2. (2)
    $M\otimes\mathbb{Z}_{p}$ splits three hyperbolic planes for all primes $p$.
Then we have